Two important industries on the island of Bermuda are fishing and tourism. According to data from the Food and Agriculture Organization of the United Nations and the Bermuda Department of Statistics, in 2014 the 315 registered fishermen in Bermuda caught 497 metric tons of marine fish. And the 2,446 people employed by hotels produced 580,209 hotel stays (measured by the number of visitor arrivals). Suppose that this production point is efficient in production. Assume also that the opportunity cost of 1 additional metric ton of fish is 2,000 hotel stays and that this opportunity cost is constant (the opportunity cost does not change).

1. If all 315 registered fishermen were to be employed by hotels (in addition to the 2,446 people already working in hotels), how many hotel stays could Bermuda produce?

2. If all 2,446 hotel employees were to become fishermen (in addition to the 315 fisherman already working in the fishing industry), how many metric tons of fish could Bermuda produce?

3. Draw a production possibility frontier for Bermuda, with fish on the horizontal axis and hotel stays on the vertical axis, and label Bermuda's actual production point for the year 2014.
Answer #1

Fish caught by one fisherman = 497 metric ton / 315 = 1.58 metric ton

Hotel stays per hotel employee = 580,209 / 2,446 = 237.21

(1)

Amount of fish given up by 315 fishermen joining a hotel = 497 metric ton

Additional hotel stays by 497 metric ton of fish given up = 497 x 2,000 = 994,000

Total number of hotel stays = 580,209 + 994,000 = 1,574,209

(2)

Number of hotel stays given up by 2,446 hotelmen becoming fishermen = 580,209

Additional amount of fish caught by 580,209 hotel stays given up = 580,209 / 2,000 = 290.10

Total amount of fish = (497 + 290.1) metric ton = 787.1 metric ton

(3)

PPF is as follows. Actual production point is at A (490 metric tons of fish and 580,209 hotel stays).

S14 209 S80 20 PPF 187. Fish (MT) 490